OpenStudy (anonymous):

If two positives equals a negative..than why is -4^2 = -16 NOT +16? Please answer thx!

OpenStudy (anonymous):

The reason is because of the order of operations. You do the exponent first, and the subtraction second.

OpenStudy (anonymous):

So we have \(-(4)(4)\) instead of \((-4)(-4)\)

OpenStudy (campbell_st):

well its simply -4 x -4 which is 2 negatives and if you are using a calculator then enter it as (-4)^2

zepdrix (zepdrix):

\[-4^2=-(4^2)\neq (-4)^2\]

OpenStudy (anonymous):

cant use a calculator :\ so its -(4x4) = -(16) = 16

OpenStudy (anonymous):

-4^2 implies -1 x 4^2 = -1 x 4 x 4 = -16 But (-4)^ 2 = -4 x -4 = 16

zepdrix (zepdrix):

amy why did your -16 magically turn into 16? :o

OpenStudy (campbell_st):

it depends on the context if its simply evaluate -4^2 then its 16 if its 3^2 - 4^2 = -7

OpenStudy (anonymous):

sorry meant -16! :)

OpenStudy (campbell_st):

you have to look at the question and decide if the sign is a subtraction operation of a negative

zepdrix (zepdrix):

it doesn't actually matter camp, eve if its an operation, it will be written very clearly if you're suppose to square the negative. like this: \[(-4)^2\] :D

OpenStudy (anonymous):

Remember that \[-4^2 \neq (-4)^2\]

OpenStudy (anonymous):

well it simply says evaluate:: -4^2 my answer was 16 and i was wrong

OpenStudy (anonymous):

> two positives equals a negative.

OpenStudy (anonymous):

THanks guys!